# Authors Should Annotate

### v1

@@ Abstract
The status quo for labeling text is third-party annotation, but there are many cases where information directly from the document’s source would be preferable over a third-person proxy, especially for egocentric features like sentiment and belief. We introduce author labeling, an annotation technique where the writer of the document itself annotates the data at the moment of creation. We collaborate with a commercial chatbot with over 10,000 users to deploy an author labeling annotation system for subjective features related to product recommendation. This system identifies task-relevant queries, generates on-the-fly labeling questions, and records authors’ answers in real time. We train and deploy an online-learning model architecture for product recommendation that continuously improves from author labeling and find it achieved a 534% increase in click-through rate compared to an industry advertising baseline running concurrently. We then compare the quality and practicality of author labeling to three traditional annotation approaches for sentiment analysis and find author labeling to be higher quality, faster to acquire, and cheaper. These findings reinforce existing literature that annotations, especially for egocentric and subjective beliefs, are significantly higher quality when labeled by the author rather than a third party. To facilitate broader scientific adoption, we release an author labeling service for the research community at academic.echollm.io.
Authors Should Annotate
Marcus Ma∗1,2   Cole Johnson∗1   Nolan Bridges∗1 Jackson Trager2Georgios Chochlakis2 Shrikanth Narayanan2 1 echo LLM  2 University of Southern California {marcus,cole,nolan}@echollm.io equal contribution
@@ 1 Introduction
The ultimate goal of language understanding is decoding the underlying intent behind a speaker’s words. Intent cannot be intuited from text alone; true understanding requires contextualizing the message with the source view shaped by the communicator’s individual preferences, beliefs, and judgments. However, when training language models to infer this context, access to the original source is often difficult or impossible. The prevailing method for interpreting source views is via observational proxy, whereby third-party annotators assign labels to text data based on their inferences of what the source author would have believed pustejovsky_2012.
Third-party observers, however, are noisy proxies of the original source. Annotators lack “privileged access” to authors’ authentic internal states vazire_2010 and lack context of the original situation jones_1971; wallace_2014. Decades of annotation research have documented fundamental limitations of third-party annotation, such as annotator biases, misalignment with the author intent, and pervasive annotator disagreement (snow_2008; fort_2016; aroyo_2015; uma_2021). While some of these limitations can be mitigated — for example, in subjective tasks, annotator disagreement may be indicative of multiple interpretations rather than noise chochlakis_2025, and standardized techniques can lead to high inter-annotator agreement paletz_2024 — they are structurally tied to the methodology of third-party annotation. The reliance on third-party annotation is simply a logistical necessity, as it is impossible to reconstruct the original author’s viewpoint during retroactive text labeling.
In this work, we investigate the technical feasibility of collecting annotation data directly from authors in real time. We present author labeling, a new annotation methodology where we present authors with the opportunity to annotate their documents at their moment of creation. In the real-world setting of chatbot conversations, we deploy a data collection system that uses lightweight LLMs to automatically monitor user messages for relevance to various topics and tasks. When this system identifies task-relevant moments, we create a contextualized annotation task based on the user’s query and present it to the user directly in the chatbot interface for them to answer.
@@
To test author labeling, we introduce Echo, or Experimental Coordination of Human Objectives, as an online-learning model architecture that continuously improves through author labeling by strategically querying authors to reduce uncertainty over a predefined set of features or tasks. We implement Echo for the task of product recommendation and deploy it into a live chatbot service with over 10,000 users simultaneously with an industry-standard banner-ad company’s product. We find that Echo achieves a click-through rate (CTR) of 7.52%, a fivefold increase over the industry baseline’s 1.40% CTR. We also directly compare our author-labeling approach to other annotation methods for the task of sentiment analysis. Using three other annotation sources — LLM-as-an-annotator, crowdsource via Mechanical Turk, and hand-selected expert — we find that author labeling is the most performant, internally consistent and cheapest (besides LLM annotation) of the methods. To support broader scientific adoption, we release a publicly available author labeling service at academic.echollm.io, enabling researchers to build datasets grounded in real-time author intent rather than third-party approximation.
@@ 2 Author Labeling
Author labeling is an analogous methodology to ecological momentary assessment (EMA), a common technique in behavioral psychology that uses egocentric, in-the-moment self-reports (shiffman_2008). In §2.1, we highlight three reasons why direct self-reports such as author labeling and EMA are preferable to third-party annotation. In §2.2, we provide an overview of how we conducted author labeling in a chatbot setting, and in §2.3, we discuss limitations of author labeling.
2.1 Motivation for Author Labeling
@@
Language is inherently ambiguous, and a single sentence can contain many different interpretations. Even the most skilled and attentive annotators can at best identify ambiguities and make an educated guess, and only the author of the document understands the true intent. oprea_2020 find that annotator performance on sarcasm detection — an example of an ambiguous linguistic construct — achieves an F1-score of 0.616 compared to authors’ actual intents. Additionally, when constructing passages, authors themselves are fully aware of the background, personal context, and other non-linguistic influences that motivate the writing of the passage in the first place. buechel_2017 find that instructions prompting annotators to explicitly consider the author’s context and viewpoint increases label quality.
@@
In subjective annotation tasks, disagreement between annotators can mean more than noise and can in some cases be a useful signal itself for detecting ambiguity beigman_2009; chochlakis_2025. This ambiguity is simply a residue of the annotation process, though; language can be ambiguous to external annotators, but to the author, the meaning should be more clear.
@@ 2.2 Data Collection
The largest obstacle for collecting data via author labeling is access to a diverse, large data source with an integrated UI platform for author annotations. We partner with echo LLM [^1] https://chat.echollm.io, a free-to-use chatbot platform that provides users with personalized product recommendations. Echo relays LLM inference to third-party LLM providers such as OpenAI and Anthropic and monitors the output stream. When recommendation for a product from echo’s catalogue would benefit the user’s conversation, echo generates a conversation-contextualized and LLM-generated advertisement prompted to be as helpful and complementary to the chatbot response as possible.
We conducted author labeling with the purpose of gathering labeled information on a variety of intent features relevant to product recommendation, such as sentiment, urgency, and desired product features. We set up an automatic trigger that ran on every user query, where after a message was sent, we first passed the contents through a spam and greetings filter. Then, we evaluated the message with a lightweight LLM to see if it a given feature was relevant to the conversation. Some features, such as sentiment, were applicable to almost every query, while others, such as purchase urgency, were only relevant when it was clear the user was exhibiting purchase intent. If enough intent features that we needed data on were deemed relevant to the conversation, we presented users with a popup overlay interface asking them four multiple-choice or single-word free-text questions relevant to the conversation (Figure 1). These tasks are generated on-the-fly based on the conversation, and each question is designed to add new information about one specific aspect of the user’s intent that is not immediately answerable from the conversation itself. We provide additional detail in §3.1.
@@
We designed our label questions to be as non-intrusive and easy-to-answer as possible, drawing on established survey design principles stating that minimizing cognitive burden reduces the likelihood of low-quality or falsified data krosnick_1991; dillman_2014; this is especially important given that the authors themselves were not presented with detailed instructions typically given to third-party annotators. With this in mind, we designed all of our questions to require either one-click or one-word answers and enforced a 5-second minimum delay after presenting the task so that users would read the question rather than rush through. We also relied on echo’s user fraud detection algorithms and only report users whose data has been verified by their system to be not fraudulent (see §B.2 for more details).
In October and November 2025, a total of 928,078 messages were sent on echo from 22,077 users. We filtered out tasks where the conversation revealed PII or discussed sensitive topics, and we also filtered out users who were spamming echo to exclusively receive tasks rather than asking genuine questions. After these filters, we presented 54,977 four-question surveys and achieved a survey completion rate of 84.3%, collecting 185,548 individual author-labeled datapoints.
@@ 3.3 Downstream Prediction
We use the features passed from the feature selector in our downstream algorithm, which is configured to take exactly $k$ input features, which has a downstream evaluation and loss. In our implementation of a product recommendation system for example, described in further detail in §4, the downstream recommendation algorithm receives specific user-intent features from the feature selector and produces the contextualized recommendation advertisement, where we use the click-through rate and more direct author labeling (i.e., asking if the recommendation was helpful) as our downstream loss.
@@ 3.4 Training
We have two forms of feedback to supervise the training of Echo: the author labeling on specific feature predictions, which is used to train the feature models and the confidence of the feature selector, and the downstream loss, which trains the feature selector’s relevance scores and optionally the downstream algorithm itself.
@@ 3.4.1 Author Labeling
After the feature selector identifies the top-$k$ features, we use the uncertainty scores $\bm{\sigma}$ and generate tasks to ask the user what the true value of that feature should be, making sure to include the predicted feature value as one of the options. When the author provides the ground-truth label for that feature, we use the value to update both the feature selector’s uncertainty calculations and the associated feature model’s prediction. We use this loss to update only the specific feature model, not all feature models, as the author labeling task is generated solely for that individual feature. To update the feature selector’s uncertainty, we binarize the uncertainty loss into a 0 or 1 depending on if the label chosen by the author matches the feature model’s predicted label.
@@ 4.2 Results
We plot the day-by-day CTR of Echo and the industry baseline in Figure 3. In total, Echo created 10,378 impressions with 789 clicks for an overall CTR of 7.52%, while the industry baseline presented 1,711 impressions with 24 clicks for a CTR of 1.40%, meaning that Echo achieves a 537% performance increase compared to the industry standard.
@@
While click-through rates for ad displays in chatbot environments have been understudied, the global average CTR for display ads on Google Display Network (i.e., on internet search) ranges between 0.46% (irvine_2024) and 0.57% (chaffey_2024), with static banner ads seeing CTR as low as 0.10% bannerflow_2024. We believe that the increased performance of the industry baseline, achieving 1.40% CTR in our chatbot environment, could be due to higher levels of attention and novelty for users using chatbots versus internet search.
@@
We display day-by-day CTR in Figure 4. We observe that Mondays and Fridays have the lowest average CTR with peaks mid-week and on the weekend, aligning with established consumer patterns. Monday’s low performance aligns with the task-orientation online browsing hypothesis Bussiere_2016, stating that consumers allocate their cognitive resources to starting the work week, so buying products is not on their mind. Mid-week jumps could be explained by “cyberloafing” Lim_2002, where online browsing increases as a distraction from work; the Friday dip could be best explained by higher rates of socialization, which is also evidenced by a dip in impressions on Friday (see Figure 9 in Appendix C), with a weekend surge in CTR possibly explained by increased leisure and exploration from online consumers Moe_2004.
@@ 5 Comparing Author Labeling to 3rd-Party Annotation
We now perform a head-to-head analysis of author labeling for annotation compared to three other annotation methods for the task of sentiment analysis. In the chatbot platform, we introduced an automatic task generator that created multiple-choice questions asking the user about their current emotional state based on their conversation. Both the question and the answers were derived from an automatic LLM generation prompted to create a contextualized question aimed at understanding the user’s emotional state and to come up with several plausible emotions or feelings as options. We intentionally chose to have the LLM come up with labels rather than select options from a predefined emotion set so that the labels themselves could incorporate context, thus further disambiguating the label value to users (i.e. the label “Concerned about pollution” vs. “Concerned”). We filtered out spam and non-English messages, messages from users detected to be bots, and messages with little to no emotional subtext. We also excluded all tasks with any mention of personally identifiable information, religion, NSFW content, non-English, or political content. We restricted our sample to users who had completed a minimum of ten annotation tasks and sampled exactly 6 tasks for every user, and performed a final manual inspection of all user messages and tasks. This yielded 1,962 total tasks across 327 users and 1043 conversations, with each conversation having an average of 1.88 tasks and each user having an average of 3.19 conversations.
@@
We report annotator demographics in Table 5; however, the authors have reason to believe that some of the demographics reported by Mechanical Turk annotators may have been falsified, as the authors received 15 emails from MTurk annotators with traditionally Indian or South Asian names written in non-fluent English, yet only two MTurk annotators reported having an Asian ethnicity.
For the expert annotators, we selected a sample of 300 sentiment tasks from the 1,962 total tasks. We collected three annotations per datum for all annotation methods, totaling 196 Mechanical Turk annotators and 30 expert annotators. For the LLM annotator, we used the same prompt for Llama 3.1 8B, Llama 3.3 70B, and GPT-OSS 20B as the three “annotators”. We present additional annotator details in Appendix B.
@@ 5.2 Alignment with Author Labels
We present overall accuracy with author labels, annotator agreement, and KL divergence with author label distributions for the three annotation methods in Table 1444We note that random accuracy is 25%, as all data are four-choice multiple-choice questions.. We find that experts (42.76%) and LLMs (42.33%) achieve essentially the same level of accuracy, but LLM annotators had much higher agreement ($\kappa$=0.36) than experts (0.12). However, we find that when experts do agree, their overall accuracy is higher (Table 2), achieving the highest accuracy among methods of 63.0% when all annotators agree on a label. Meanwhile, MTurk annotations perform the worst on all three metrics, which is to be expected compared to expert annotators but surprising in comparison to LLM annotation.

### v2

# Authors Should Label Their Own Documents

@@ Abstract
Third-party annotation is the status quo for labeling text, but egocentric information such as sentiment and belief can at best only be approximated by a third-person proxy. We introduce author labeling, an annotation technique where the writer of the document itself annotates the data at the moment of creation. We collaborate with a commercial chatbot with over 20,000 users to deploy an author labeling annotation system. This system identifies task-relevant queries, generates on-the-fly labeling questions, and records authors’ answers in real time. We train and deploy an online-learning model architecture for product recommendation with author-labeled data to improve performance. We train our model to minimize the prediction error on questions generated for a set of predetermined subjective beliefs using author-labeled responses. Our model achieves a 537% improvement in click-through rate compared to an industry advertising baseline running concurrently. We then compare the quality and practicality of author labeling to three traditional annotation approaches for sentiment analysis and find author labeling to be higher quality, faster to acquire, and cheaper. These findings reinforce existing literature that annotations, especially for egocentric and subjective beliefs, are significantly higher quality when labeled by the author rather than a third party. To facilitate broader scientific adoption, we release an author labeling service for the research community at academic.echollm.io.
Authors Should Label Their Own Documents
Marcus Ma∗1,2   Cole Johnson∗1   Nolan Bridges∗1 Jackson Trager2Georgios Chochlakis2 Shrikanth Narayanan2 1 Echo Group  2 University of Southern California {marcus,cole,nolan}@echollm.io equal contribution
@@ 1 Introduction
Text alone cannot intuit intent. True language understanding requires contextualizing messages with communicators’ source views shaped by individual preferences, beliefs, and judgments. However, when training language models to infer this context, access to the original source is often difficult or impossible. The prevailing method for interpreting source views is via observational proxy, whereby third-party annotators assign labels to text data based on their inferences of what the source author would have believed pustejovsky_2012.
Third-party observers, however, are noisy proxies of the original source. Annotators lack “privileged access” vazire_2010 to authors’ authentic internal states, meaning they lack situational context that can cause misinterpretations of original intent jones_1971; wallace_2014. Decades of annotation research have documented fundamental limitations of third-party annotation, such as annotator biases, intent misalignment, and pervasive annotator disagreement (snow_2008; fort_2016; aroyo_2015; uma_2021). While some of these limitations can be mitigated, such as through inter-annotator analysis paletz_2024; chochlakis_2025, these problems are structurally tied to third-party annotation. Reliance on third-party annotation is simply a logistical necessity, as it is impossible to reconstruct the original viewpoint during retroactive text labeling.
In this work, we investigate the technical feasibility of collecting annotation data directly from authors in real time. We present author labeling, a new annotation methodology where we present authors with the opportunity to annotate their documents at their moment of creation with LLM-generated questions. In the real-world setting of chatbot conversations, we deploy a data collection system that uses lightweight LLMs to automatically monitor user messages for relevance to various topics and tasks. When this system identifies task-relevant moments, we create a contextualized annotation task based on the user’s query and present it to the user directly in the chatbot interface for them to answer.
@@
To test author labeling, we introduce Echo, or Experimental Coordination of Human Objectives, an online-learning model architecture that continuously improves through author labeling by strategically querying authors to reduce uncertainty over a predefined set of features or tasks. We implement Echo and deploy it into a live chatbot service with over 10,000 users for the task of product recommendation, a task that strongly depends on individual sentiment and preference with an objective measure of success, i.e. a click on the presented advertisement. We ran Echo side-by-side with an industry-standard banner-ad company’s product and find Echo achieves a click-through rate (CTR) of 7.52%, a fivefold increase over the industry baseline’s 1.40% CTR. We also directly compare our author-labeling approach to other annotation methods for the task of sentiment analysis. Using three other annotation sources — LLM-as-an-annotator, crowdsource via Mechanical Turk, and hand-selected expert — we find author labeling is the most performant and internally consistent, and cheapest of the human methods. To support scientific adoption of author labeling, we release a public author labeling service at academic.echollm.io, enabling researchers to build datasets grounded in real-time author intent instead of third-party approximation.
@@ 2 Author Labeling
Author labeling provides the same theoretical benefits as other methods utilizing egocentric, real-time assessments such as patient-reported outcomes patrick_2011_content; FDA_PRO_2009 in the medical field and ecological momentary assessment (shiffman_2008) in psychology. In §2.1, we highlight three advantages of author labeling over third-party annotation. In §2.2, we provide an overview of how we conducted author labeling in a chatbot setting, and in §2.3, we discuss limitations of author labeling.
2.1 Advantages of Author Labeling
@@
Language is inherently ambiguous and can contain many valid interpretations. Skilled annotators can identify these ambiguities, but they have no means of resolving them. oprea_2020 find annotator performance on sarcasm detection — an example of an ambiguous linguistic construct — achieves an F1-score of 0.616 compared to authors’ actual intents. Annotators often fail to consider the full background, personal context, and other non-linguistic influences that motivate the writing of the passage; buechel_2017 find instructions prompting annotators to explicitly consider the author’s context and viewpoint increases label quality.
@@
In subjective annotation tasks, disagreement between annotators can mean more than noise and can in some cases be a useful signal itself for detecting ambiguity beigman_2009; chochlakis_2025. This ambiguity is simply a residue of the annotation process, though; language can be ambiguous to external annotators, but to the author, the meaning should be more clear [^1] Whether meaning is clear to the original speaker is itself a contested belief nisbett_1977. To this we assert meaning is at least more apparent to the author than to a proxy..
@@ 2.2 Data Collection
The largest obstacle for collecting data via author labeling is access to a diverse, large data source with an integrated UI platform for author annotations. We partner with Echo Group [^2] https://chat.echollm.io, a free-to-use chatbot platform that provides users with personalized product recommendations. Echo Group relays LLM inference to third-party LLM providers such as OpenAI and Anthropic and monitors the output stream. When recommendation for a product from Echo Group’s catalog would benefit the user’s conversation, it generates a conversation-contextualized and LLM-generated advertisement prompted to be as helpful and complementary to the chatbot response as possible.
We conducted author labeling with the purpose of gathering labeled information on a variety of intent features relevant to product recommendation, such as sentiment, urgency, and desired product features. We set up an automatic trigger that ran on every user query, where after a message was sent, we first passed the contents through a spam and greetings filter. We then evaluated the message with a lightweight LLM to see if it a given feature was relevant to the conversation. Some features, such as sentiment, were applicable to almost every query, while others, such as purchase urgency, were only relevant when the user was clearly exhibiting purchase intent. If enough intent features were deemed relevant to the user message, we presented users with a popup overlay asking them four multiple-choice or single-word free-text questions (Figure 1). These tasks are contextually generated on-the-fly, with each question designed to gain new information about its associated feature. We provide additional detail in §3.1.
@@
We designed our label questions to be as non-intrusive and easy-to-answer as possible, drawing on established survey design principles that minimizing cognitive burden reduces the likelihood of low-quality or falsified data krosnick_1991; dillman_2014; this is especially important given authors are not presented with detailed instructions typically given to third-party annotators. We designed all our questions to require either one-click or one-word answers and enforced a 5-second minimum reading time after presentation. We also relied on Echo Group’s user fraud detection algorithms and only report data from trustworthy users (see §B.2).
In October and November 2025, we processed 928,078 messages from 22,077 users. After filtering PII, sensitive topics, and spam, we displayed 54,977 four-question surveys and recieved 185,548 individual author-labeled datapoints, achieving a survey completion rate of 84.3%.
@@ 3.3 Downstream Prediction
We use the features passed from the feature selector in our downstream algorithm, which is configured to take exactly $k$ input features, which has a downstream evaluation and loss. In our implementation of a product recommendation system for example, described in further detail in §4, the downstream recommendation algorithm receives specific user-intent features from the feature selector and produces the contextualized recommendation advertisement, where we use the click-through rate and more author labeling feedback (i.e., directly asking if the recommendation was helpful) as our downstream loss.
@@ 3.4 Training
We have two forms of feedback to supervise the training of Echo: the author labeling on specific feature predictions, which is used to train the feature models and the confidence of the feature selector; and the downstream loss, which trains the feature selector’s relevance scores and optionally the downstream algorithm itself.
@@ 3.4.1 Author Labeling
After the feature selector identifies the top-$k$ features, we use the uncertainty scores $\bm{\sigma}$ and generate tasks to ask the user what the true value of that feature should be, making sure to include the predicted feature value as one of the options. When the author provides the ground-truth label for that feature, we use the value to update both the feature selector’s uncertainty calculations and the associated feature model’s prediction. We use this loss to update only the specific feature model, not all feature models, as the author labeling task is generated solely for that individual feature. To update the feature selector’s uncertainty, we binarize the uncertainty loss into a 0 or 1 depending on if the label chosen by the author matches the feature model’s predicted label. For free-response, loss is the cosine distance between predicted and author-labeled text embeddings.
@@ 4.2 Results
We plot the day-by-day CTR of Echo and the industry baseline in Figure 3. In total, Echo created 10,378 impressions with 789 clicks for an overall CTR of 7.52%, while the industry baseline presented 1,711 impressions with 24 clicks for a CTR of 1.40%, a 537% performance increase.
@@
While click-through rates for ad displays in chatbot environments have been understudied, the global average CTR for display ads on Google Display Network (i.e., on internet search) ranges between 0.46% (irvine_2024) and 0.57% (chaffey_2024), with static banner ads seeing CTR as low as 0.10% bannerflow_2024. We believe the industry baseline’s CTR of 1.40% could be due to higher levels of attention and novelty for users using chatbots versus internet search.
@@
We display day-by-day CTR in Figure 4. We observe Mondays and Fridays have the lowest average CTR with peaks mid-week and on the weekend, aligning with established consumer patterns. Monday’s low performance aligns with the task-orientation online browsing hypothesis Bussiere_2016, stating consumers allocate their cognitive resources to starting the work week, so buying products is not on their mind. Mid-week jumps could be explained by “cyberloafing” Lim_2002, where online browsing increases as a distraction from work; the Friday dip could be best explained by higher rates of socialization, which is also evidenced by a dip in impressions on Friday (see Figure 9 in Appendix C), with a weekend surge in CTR possibly explained by increased leisure and exploration from online consumers Moe_2004.
@@ 5 Comparing Author Labeling to 3rd-Party Annotation
We also perform a head-to-head analysis of author labeling for annotation compared to three status quo annotation methods. We use Echo’s multiple-choice questions generated for sentiment, where an LLM was probed to understand the user’s emotional state and present four plausible emotions as options. We opted to use LLM-generated label options rather than predefined emotion labels so that the labels themselves could incorporate context for maximum user clarity (i.e. the label “Concerned about pollution” vs. “Concerned”). We apply the same filtering criteria as §2.2 and restricted our sample to users who had completed a minimum of ten annotation tasks and sampled exactly six tasks for every user, applying a final manual inspection over all user messages and tasks. This yielded 1,962 total tasks across 327 users and 1043 conversations, with an average of 1.88 tasks per conversation and 3.19 conversations per user.
@@
We report annotator demographics in Table 5; however, the authors have reason to believe that some of the demographics reported by Mechanical Turk annotators may have been falsified, as the authors received 15 emails from MTurk annotators with traditionally Indian or South Asian names written in non-fluent English yet only two MTurk annotators reported having an Asian ethnicity.
For the expert annotators, we selected a sample of 300 tasks from the 1,962 total. For LLM annotators, we used the same prompt for Llama 3.1 8B, Llama 3.3 70B, and GPT-OSS 20B as the three “annotators”. We collected three annotations per datum for all annotation methods using 196 Mechanical Turk annotators, 30 expert annotators, and 5,886 LLM calls. We present additional details and prompts in Appendix B.
@@ 5.2 Alignment with Author Labels
We present overall accuracy with author labels, annotator agreement, and KL divergence with author label distributions for the three annotation methods in Table 1555We note that random accuracy is 25%, as all data are four-choice multiple-choice questions.. We find experts (42.76%) and LLMs (42.33%) achieve essentially the same level of accuracy, but LLM annotators had much higher agreement ($\kappa$=0.36) than experts (0.12). However, we find that when experts do agree, their overall accuracy is higher (Table 2), achieving the highest accuracy among methods of 63.0% when all annotators agree on a label. Meanwhile, MTurk annotations perform the worst on all three metrics, which is to be expected compared to expert annotators but surprising in comparison to LLM annotation.